OpenStudy (anonymous):

what is the probability that a number selected at random form the first 1,000 positive integers is exactly divisible by 6 or 9?

OpenStudy (anonymous):

How many numbers in that range are divisble by 6? \[ \lfloor\frac{1000}{6}\rfloor \] How many numbers in that range are divisble by 9? \[ \lfloor\frac{1000}{9}\rfloor \] How many numbers are divisble by 6 AND 9 ? \[ \lfloor\frac{1000}{18} \rfloor \] (Note that 18 is the lcm of 6 and 9.) So how many numbers in the range are disible by 6 OR 9? \[ \lfloor\frac{1000}{6}\rfloor+\lfloor\frac{1000}{9}\rfloor-\lfloor\frac{1000}{18} \rfloor \]

OpenStudy (anonymous):

That evaluates to 222. So your probability is \[ \frac{222}{1000} = 22.2 \% \]

OpenStudy (anonymous):

how did you get 222?

OpenStudy (anonymous):

Evaluate the expressions above.

OpenStudy (anonymous):

ook i see now.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hit good answer to let ppl know this one is solved